OpenStudy (anonymous):

Find domain of the function 1/(x+|x|)

OpenStudy (anonymous):

\[\frac{ 1 }{ x+|x|}\]

OpenStudy (zzr0ck3r):

the only problem would be when x+|x| = 0 this is true for all x <= 0 so what is the domain?

OpenStudy (zzr0ck3r):

@Palmo4ka ?

OpenStudy (anonymous):

\[(-\infty;+\infty)\] if \[x \neq-1\] ?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

do you not see me typing?

OpenStudy (zzr0ck3r):

x cant be 0, or negative. put in any negative number or 0 and you get 1/0

OpenStudy (zzr0ck3r):

no

OpenStudy (zzr0ck3r):

1/(-7+|-7|) = 1/0

OpenStudy (anonymous):

(0,inf)

OpenStudy (zzr0ck3r):

not 0 either

OpenStudy (zzr0ck3r):

as 1/(0+|0|) = 1/0

OpenStudy (anonymous):

( means not included , [ means included

OpenStudy (anonymous):

all values ​​except x = -1

OpenStudy (anonymous):

? @zzr0ck3r

OpenStudy (anonymous):

From 0 to positive infinity, 0 is NOT included

OpenStudy (zzr0ck3r):

yes word, sorry its late

OpenStudy (anonymous):

okay, i understood, thank you

OpenStudy (zzr0ck3r):

d := {x|x>0}

OpenStudy (anonymous):

Help me with else one question, plz \[y=\frac{ 1 }{ x-|x| }\] The domain is \[ (-\infty ; 0)\] ?

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

dont use (a;b) use (a,b)

OpenStudy (anonymous):

ok! Thanks!

OpenStudy (zzr0ck3r):

np